OpenStudy (anonymous):

find the product. simplify if possible 5 over 14 times 4a over 25 A. a over 35 B. 2 over 35 C. 2a over 35 D. a over 30

OpenStudy (jdoe0001):

\(\bf \cfrac{5}{14}\times \cfrac{41}{25}\implies \cfrac{5\times 4a}{14\times 25}\)

OpenStudy (anonymous):

so is that C @jdoe0001 ?

OpenStudy (jdoe0001):

well. you said you got C, so, is it?

OpenStudy (anonymous):

idk, that was for a different question

OpenStudy (jdoe0001):

hm ok... well.. so.. multiply as then simplify

OpenStudy (anonymous):

i don't know how :/

OpenStudy (jdoe0001):

well... what would you get, without simplifying just multiplying, for \(\bf \cfrac{5\times 4a}{14\times 25}\quad ?\)

OpenStudy (anonymous):

20 over 350

OpenStudy (jdoe0001):

so \(\bf \cfrac{5}{14}\times \cfrac{41}{25}\implies \cfrac{5\times 4a}{14\times 25}\implies \cfrac{20a}{350}\qquad then\)  \(\bf \cfrac{20a}{350}\implies \cfrac{\cancel{2}\cdot 2\cdot \cancel{5}a}{\cancel{2}\cdot \cancel{5}\cdot 5\cdot 7}\)

OpenStudy (anonymous):

ok, so A or D?

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{2}\cdot 2\cdot \cancel{5}a}{\cancel{2}\cdot \cancel{5}\cdot 5\cdot 7} \implies ?\)

OpenStudy (jdoe0001):

what's left?

OpenStudy (anonymous):

A/35

OpenStudy (anonymous):

it is A

OpenStudy (jdoe0001):

notice, we only could cancel out 1 of the "2" above

OpenStudy (anonymous):

right

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{2}\cdot 2\cdot \cancel{5}a}{\cancel{2}\cdot \cancel{5}\cdot 5\cdot 7} \implies \cfrac{2a}{35}\)

OpenStudy (anonymous):

so it was C!!!!

OpenStudy (jdoe0001):

yes, but the idea is to multiply and then simplify, like so